OpenStudy (anonymous):

Please help.. wite complex number in standard form: \[\frac{ 5+4i }{ -9i }\]

OpenStudy (anonymous):

If you have a complex number: \[z=x+iy\] With x and y real then you define division by complex numbers as multiplying top and bottom by the complex conjugate defined as: \[z^*=x-iy\] Please also note that: \[zz^*=x^2+y^2\] Therefore division is defined as: \[\frac{w}{z}=\frac{w z^*}{z z^*}=\frac{w z^*}{x^2+y^2}\] So we have:\[\frac{5+4i}{-9i}=\frac{(5+4i)(+9i)}{(-9i)(+9i)}=\frac{45i-36}{81}=\frac{5i-4}{9}\]

OpenStudy (anonymous):

Where if we let z=x+iy then x=0 for the denominator.